# Adapting Speech Foundation Models with Large Language Models for Unified Speech Recognition

### v1

@@ Abstract
Unified speech recognition aims to perform auditory, visual, and audiovisual speech recognition within a single model framework. While speech foundation models (SFMs) have demonstrated remarkable performance in auditory tasks, their adaptation to multimodal scenarios remains underexplored. This paper presents UASR-LLM, a novel framework that adapts frozen SFMs to unified visual speech recognition (VSR), automatic speech recognition (ASR), and audio-visual speech recognition (AVSR) tasks by leveraging large language models (LLMs) as text decoders. Our approach introduces visual representations into multiple SFM layers through visual injection modules, enabling multimodal input processing and unified hidden representations. The augmented SFMs connect with decoder-only LLMs via a feed-forward adaptor, where concatenated representations and instruction prompts guide speech transcription. We propose a two-stage training strategy: visual injection pretraining followed by speech recognition finetuning. SFM parameters remain frozen throughout training, with only visual injection modules optimized initially, and LLMs finetuned using LoRA parameters subsequently. Experimental results demonstrate superior performance over state-of-the-art baselines across VSR, ASR, and AVSR tasks under both clean and noisy conditions. Ablation studies confirm generalization across various SFMs and LLMs, validating the proposed training strategy.
keywords:
audio-visual speech recognition , unified speech recogntion , speech foundation model , lipreading , large language models[snnu]organization=School of Artificial Intelligence and Computer Science, Shaanxi Normal University,addressline=No. 620, West Chang’an Avenue, city=Xi’an, postcode=710119, state=Shaanxi, country=P.R.China
[ustc]organization=University of Science and Technology of China,addressline=No.96, JinZhai Road, city=Hefei, postcode=230026, state=Anhui, country=P.R.China
[ifly]organization=iFLYTEK Research, iFLYTEK Co. Ltd.,addressline=No. 666, Wangjiang West Road, city=Hefei, postcode=230088, state=Anhui, country=P.R.China
1 Introduction
Speech recognition represents a fundamental challenge in pattern recognition [1], with recent advances driven by speech foundation models (SFMs) such as WavLM [2] and Whisper [3]. These large-scale models, trained on extensive speech datasets, have demonstrated remarkable performance in automatic speech recognition (ASR). However, environmental noise remains a critical limitation that can significantly compromise recognition accuracy, particularly in real-world deployment scenarios. To overcome this challenge, researchers have developed complementary approaches including visual speech recognition (VSR, also known as lipreading) [4, 5, 6] and audio-visual speech recognition (AVSR) [7, 8], which leverage lip movement information either independently or in combination with audio signals. AVSR systems demonstrate substantial improvements in robustness under noisy conditions, making them increasingly valuable for practical applications such as intelligent home systems, online conferencing, and service robotics.
AVSR models typically adopt a sequence-to-sequence architecture nowadays, including recurrent neural network transducer (RNN-T) [9] or hybrid connectionist temporal classification (CTC) and attention architectures[10, 11]. Recently, audio-visual self-supervised learning has also been proposed to leverage unlabeled audio-visual data for pretraining, such as AV-HuBERT [12], REVAn [13], and AV2vec [13]. With a further finetuning stage on few labeled data, these methods achieve a strong VSR and AVSR performance. Building on these progresses, Haliassos et al. introduced unified speech recognition (USR) [14], which consolidates ASR, VSR, and AVSR tasks within a single model framework. This unified approach offers compelling advantages by eliminating the need to construct separate models for each modality, thereby significantly reducing training and deployment costs. The conceptual foundation is particularly appealing, as audio and visual lip movements represent complementary signal views of the same articulatory process during speech production.
Recent works have explored utilizing SFMs for VSR [15, 16] and AVSR [17, 9]. Our previous study [18] proposed using SFMs for knowledge distillation in audio-visual self-supervised pretraining, achieving competitive performance across VSR, ASR, and AVSR benchmarks. However, the knowledge distillation approach may inherently limit performance compared to direct utilization of SFMs, particularly in high signal-to-noise ratio (SNR) scenarios where the original models excel. Several approaches have focused on the direct utilization of SFMs for audio-visual tasks. Lip2vec [15] proposed reprogramming a wav2vec 2.0 model for lipreading by feeding transformed visual representations to the audio model. Whisper-flamingo [19] adapted the Whisper model for AVSR by incorporating an additional visual encoder to process visual information alongside audio inputs.
Despite its success, most previous works address VSR, AVSR, and ASR tasks separately. In this study, we explore tackling these tasks within a single unified model. Furthermore, we aim to leverage the rich knowledge embedded in both speech foundation models (SFMs) and large language models (LLMs) to design a general framework that achieves robust USR performance. We propose UASR-LLM (Unified Adaptive Speech Recognition with Large Language Models), a novel framework that adapts frozen speech foundation models to unified VSR, ASR, and AVSR tasks by leveraging LLMs as text decoders. Our approach employs a visual encoder from our previous study [18] to extract visual representations, which are then integrated into each block of the SFMs through visual injection modules. The SFM outputs are processed by an adapter to obtain hidden embeddings, which are subsequently fed into LLMs for conducting speech recognition across all modalities. To enable SFMs to process visual and audio-visual inputs, we introduce a visual injection pretraining stage based on cross-modal knowledge distillation loss. The model then undergoes a second training stage for speech recognition, incorporating random modality dropout guided by corresponding instructions. Throughout the entire training process, both the SFM and visual encoder parameters remain frozen, while the LLM is finetuned using a small set of Low-Rank Adaptation (LoRA) weights [20]. This approach enables memory-efficient finetuning while preserving the generalization capabilities of the original models. By utilizing the SFM as a unified backbone for processing multimodal audio and visual inputs, our framework facilitates knowledge transfer from audio to video data and unifies the hidden representations of both modalities. This provides a well-aligned output space for LLMs to generate transcriptions, seamlessly achieving VSR, ASR, and AVSR within a single model.
@@ 1 Introduction
Our experimental evaluation demonstrates that UASR-LLM consistently outperforms state-of-the-art baselines across all three tasks when trained with comparable amounts of audio-visual speech data. On the LRS3 test set, UASR-LLM achieves word error rates of 20.9%, 0.84%, and 0.69% for VSR, ASR, and AVSR, respectively. Under noisy conditions, our method exhibits substantial robustness, significantly reducing WER compared to baseline approaches. Furthermore, we demonstrate that our approach generalizes effectively across different speech foundation models and large language models, with all configurations outperforming baseline methods. Finally, we validate our training strategy and confirm the effectiveness of performing unified speech recognition within a single model architecture.
2 Related works
2.1 Speech foundation models
@@
Recent years have witnessed the emergence of various speech foundation models, which can be broadly categorized into two paradigms: self-supervised and supervised approaches. Among SSL-based speech foundation models, HuBERT [21] utilizes k-means clustering to generate pseudo-class labels for model training using a mask prediction loss, with iterative refinement through subsequent clustering and mask prediction steps. WavLM [21] improves HuBERT by using a more diverse pretraining dataset and performs speech denoising modeling during pretraining. On the other hand, Whisper family [3] exemplifies supervised learning-based speech foundation model. Trained on large-scale multilingual and multitask labeled data, it demonstrates robust performance in ASR and speech translation tasks.
2.2 Visual representation learning
@@
Visual representation learning is crucial for the performance of VSR and AVSR tasks, which can be broadly categorized into supervised and self-supervised approaches. Supervised methods leverage labeled [22] or weakly labeled audio-visual speech data [23] to optimize representation extraction and text decoding in an end-to-end manner through label prediction. Several auxiliary tasks have been proposed to enhance visual encoder learning, including articulation feature prediction [24] and distillation from acoustic models [24, 25, 26]. Self-supervised methods utilize unlabeled audio-visual data with pretext tasks such as masked prediction loss functions. Representative approaches include AV-HuBERT [12], RAVen [13, 27], AV2vec [28], and AV-data2vec [29]. Our previous work proposed a joint audio-visual representation learning method [18] that leverages cross-modal knowledge distillation from pretrained SFMs. This work adopts our previously pretrained model for visual representation extraction, which can be viewed as an extension of our previous study. Here, we further leverage SFMs directly for processing audio-visual data, moving beyond the distillation approach to direct utilization of foundation models.
2.3 Speech recognition with LLMs
@@
Large language models (LLMs) have attracted significant attention for speech recognition applications, with research broadly falling into two categories: multimodal LLMs for universal AI assistants [30, 31] and LLMs for enhancing traditional speech recognition tasks [32]. Our work belongs to the latter category. Recent approaches for integrating LLMs with speech recognition typically employ lightweight adapters to connect frozen speech encoders with LLMs. Representative works include joint speech-language models with small adapters [32] and Q-Former-based connections [33]. Some studies explore discrete speech representations as LLM inputs [30, 34]. LLMs have also been adopted for visual and audio-visual speech recognition. VSP-LLM [35] focuses on visual speech recognition, while Llama-AVSR [36] uses separate audio and visual encoders with temporal concatenation as LLM input. MMS-Llama [37] employs audio-visual fusion within encoders and Q-former adapters for improved efficiency. Our work differs by proposing a unified speech recognition (USR) approach that processes both audio and visual modalities within a single model architecture, rather than targeting specific modalities separately.
3 Proposed method
@@
Our UASR-LLM method comprises two main components: an audio-visual adapted speech foundation model (SFM-AV) that processes both audio and visual modalities and encodes them into a unified representational space, and a large language model that decodes the encoded representations into text. The SFM-AV consists of a visual encoder and a speech foundation model enhanced with visual injection modules. Our model architecture is illustrated in Figure 1. The training of our model follows a two-stage process: in the first pretraining stage, the SFM is adapted to encode both audio and visual information into a unified hidden space; in the second finetuning stage, the adapted SFM is connected to the LLM to enable unified speech recognition. The detailed model architecture and training strategy are described in the following sections.
3.1 Model structure
3.1.1 Visual encoder
@@
Let lip video be , where and represent the temporal, height, width, and channel dimensions, respectively. The visual encoder processes the visual input into hidden representations: , where is the feature dimension of the encoded representations. In this work, we adopt the visual component of the pretrained DistillAV model [18] as our visual encoder. DistillAV is designed to distill knowledge from speech foundation models (SFMs) for joint audio-visual representation learning. Unlike SFMs, which are trained on large-scale speech-only corpora, DistillAV is trained on relatively smaller audio-visual datasets. As a result, its performance in audio-only tasks is inferior to that of its teacher models, as shown in our previous ASR evaluations. However, DistillAV exhibits strong visual modeling capabilities and consistently outperforms state-of-the-art lipreading models such as AV-HuBERT [12] and AV2vec [28]. Therefore, we use the pretrained DistillAV model to extract representations from lip video inputs.
3.1.2 Visual injection module
@@ 3.1.3 Prompting the LLM for USR
3.2 Training strategy
3.2.1 Visual injection pretraining
@@
where represents instance normalization. This multi-layer aggregation captures rich hierarchical features, which follows our previous study [18].
@@
where represents learnable projection parameters. This first-stage training leverages unlabeled audio-visual speech data for unsupervised pretraining. And the large language model are not involved at this stage.
3.2.2 Speech recognition finetuning
@@
4 Experiments and results
4.1 Implementation details
@@
Model configurations. Our experiments employed a default configuration consisting of three main components: the DistillAV model as the visual encoder, WavLM-large [2] (317M parameters) as the speech foundation model (SFM), and Qwen 2.5-7B555https://huggingface.co/Qwen/Qwen2.5-7B as the large language model. The visual encoder offered two variants: Base (103M parameters) and Large (325M parameters). Following our previous study, we denote the proposed method as UASR-LLM-B and UASR-LLM-L accordingly. The visual injection modules were configured with 8 heads, 512 hidden dimensions, and 256 dimensions for both cross-attention and feed-forward networks, totaling approximately 58M parameters. The two-layer feed-forward network adapter that connects the SFM and LLMs contains around 10M parameters. LoRA weights were configured with a rank of 16, resulting in 12M total parameters. To demonstrate the generalization capability of our method, we explored additional speech foundation models: the open-source Whisper-medium [3] and an internal iFLYTEK-Speech model, both containing approximately 300M parameters comparable to WavLM-large. We also evaluated ChatGLM v3-6B666https://huggingface.co/THUDM/chatglm3-6b as an alternative large language model.
The WavLM model generates representations at 50 fps, double the frame rate of visual representations. We set relative positional encoding according to the real-time interval between audio and visual frames using sinusoidal positional embedding. The parameter in Equation 4 was set to 8 for WavLM-based SFM and 1 for Whisper and iFLYTEK-Speech-based SFM, following our previous study [18].
@@
4.2 Comparison with state-of-the-art baselines
| Methods | Unlab. | Lab. | Shared | VSR | ASR | AVSR | ASR (N) | AVSR (N) |
@@ 4.2 Comparison with state-of-the-art baselines
| Methods | Unlab. | Lab. | Shared | VSR | ASR | AVSR | ASR (N) | AVSR (N) |
@@
Table 1 and Table 2 present experimental results for models finetuned with low-resource (30h) and high-resource (433h or 1759h) labeled data, respectively. The unlabeled data refers to audio-visual data used for pretraining the DistillAV-based visual encoder and visual injection pretraining, while the labeled data denotes audio-visual data employed for speech recognition finetuning.
As shown in Table 1, our method outperforms most baselines across VSR, ASR, and AVSR tasks under both clean and noisy conditions. Although underperforming Llama-AVSR in ASR, our approach provides a unified framework effective across all three tasks. The superior VSR performance particularly demonstrates the effectiveness of ”reprogramming” WavLM for lipreading with LLMs, consistent with previous studies [15]. With 433h labeled data, as shown in Table 2, our method achieves significant accuracy improvements. Compared to previous studies using comparable audio-visual data, our approach demonstrates superior performance across all tasks under both clean and noisy conditions. With 1759h labeled data, our method achieves WERs of 20.9%, 0.84%, and 0.69% on the clean LRS3 test set, rivaling or surpassing state-of-the-art methods. However, our VSR results still lag behind LP-Conformer [43], which uses approximately 100k hours of audio-visual data.
@@
4.3 AVSR results on noisy test sets
To evaluate robustness under noisy conditions, we constructed four test sets by mixing the clean audio from the LRS3 test set with different noise types: speech, babble, natural, and music, as described in Section 4.1. We compared our approach against multiple baselines, including two audio-only SFM models (WavLM and iFLYTEK-Speech) and three AVSR models (AV-HuBERT [12], AV2vec-MLM [28], and DistillAV [18]). All audio-only models were finetuned on the LRS3 dataset. The experiments utilized 433h data for pretraining, with all models adopting the Base configuration. We used either 30h or 433h of LRS3 data for finetuning. The experimental results are summarized in Figure 4.
Our proposed UASR-LLM method (red line) consistently outperforms all baselines across different noise types and SNR levels, with particularly pronounced advantages in severe noise conditions below 0 dB. Among the four noise conditions, babble noise proves most challenging, with all methods exhibiting higher WER at low SNRs, particularly below 0 dB. However, the proposed approach maintains substantially lower error rates than competing methods even in these challenging low-SNR scenarios. The results highlight the effectiveness of incorporating visual information, as audio-only models (WavLM and iFLYTEK-Speech) degrade rapidly in noisy conditions compared to AVSR approaches.
| Lab. | LLM | SFM | VSR | ASR | AVSR | ASR (N) | AVSR (N) |
| DistillAV | |||||||

### v2

# Adapting Speech Foundation Models for Unified Multimodal Speech Recognition with Large Language Models

@@ Abstract
While speech foundation models (SFMs) have demonstrated remarkable performance in audio-only tasks, their adaptation to multimodal scenarios remains underexplored. This work presents UASR-LLM, a novel framework that adapts frozen SFMs to unified visual speech recognition (VSR), automatic speech recognition (ASR), and audio-visual speech recognition (AVSR) by leveraging large language models (LLMs) as text decoders. Visual representations are injected into multiple SFM layers via visual injection modules, enabling multimodal fusion and unified representation learning. The augmented SFMs are connected to decoder-only LLMs through a feed-forward adaptor, where concatenated representations and instruction prompts guide transcription. We propose a two-stage training strategy consisting of visual injection pretraining followed by speech recognition finetuning. The pretraining stage aligns audio, visual, and audio-visual representations within the frozen SFM backbone, while the finetuning stage integrates LLMs for unified optimization across speech recognition tasks. Experimental results demonstrate superior performance over state-of-the-art baselines across VSR, ASR, and AVSR under both clean and noisy conditions. Ablation studies further confirm generalization across various SFMs and LLMs, validating the effectiveness of the proposed training strategy.
I Introduction
Speech recognition [1, 2, 3] has advanced substantially with the emergence of speech foundation models (SFMs), such as WavLM [4] and Whisper [5], which achieve strong performance through large-scale pretraining. Nevertheless, recognition accuracy remains vulnerable to environmental noise. To improve robustness, visual speech recognition (VSR) [6, 7, 8] and audio-visual speech recognition (AVSR) [9, 10, 11] have been widely explored, exploiting lip movements either as an alternative or as a complement to audio signals. By integrating visual and acoustic cues, AVSR systems substantially enhance noise robustness and have shown increasing practical value in scenarios such as online conferencing and embodied agents.
Modern AVSR systems predominantly adopt sequence-to-sequence architectures, including recurrent neural network transducers (RNN-T) [12] and hybrid CTC–attention frameworks [13, 14, 9]. In parallel, audio-visual self-supervised learning [6, 15, 16] has emerged as an effective pretraining paradigm for leveraging large-scale unlabeled data, with representative models such as AV-HuBERT [15, 17], REVAn [18, 16], and AV2vec [18]. After finetuning on limited labeled data, these models achieve strong performance on both VSR and AVSR tasks. Building on these advances, Haliassos et al. proposed unified speech recognition (USR) [19], which jointly supports ASR, VSR, and AVSR within a single model, substantially reducing training and deployment costs compared to modality-specific systems. More recently, speech foundation models (SFMs) have been explored for VSR [20, 21] and AVSR [22, 12]. Zhang et al. [23] further demonstrated that SFMs can serve as effective teachers for knowledge distillation in audio-visual self-supervised pretraining, achieving competitive results across ASR, VSR, and AVSR benchmarks. In addition, several studies have investigated the direct adaptation of SFMs to audio-visual settings, such as Lip2vec [20] and Whisper-Flamingo [24, 25].
Despite recent progress, most existing approaches still treat ASR, VSR, and AVSR as separate tasks with modality-specific architectures. While USR [19] builds a unified model from scratch, we instead adapt existing large-scale pretrained models for unified multimodal speech recognition in a parameter-efficient manner. To this end, we propose UASR-LLM (Unified Adaptive Speech Recognition with Large Language Models), which adapts frozen speech foundation models for unified ASR, VSR, and AVSR by employing an LLM as a universal text decoder. Specifically, visual representations extracted from a pretrained visual encoder are injected into each SFM block via lightweight visual injection modules, enabling cross-modal integration while preserving the pretrained acoustic structure. The resulting multimodal SFM outputs are projected through a learnable adapter into the hidden embedding space of the LLM, which directly performs text generation across all input modalities, thereby establishing a unified decoding framework.
To enable SFMs to process visual and audio-visual inputs, we introduce a visual injection pretraining stage based on cross-modal knowledge distillation. This is followed by a unified speech recognition finetuning stage, where instruction-guided random modality dropout is applied to enhance robustness. Throughout training, both the SFM and visual encoder remain frozen, while the LLM is efficiently finetuned using low-rank adaptation (LoRA) [26]. This design substantially reduces memory and computational costs while preserving the generalization ability of pretrained models. By using a single SFM as a shared multimodal backbone, our framework facilitates knowledge transfer from audio to visual inputs and aligns representations across modalities, enabling the LLM to generate transcriptions in a unified output space and seamlessly perform ASR, VSR, and AVSR within one model.
@@ I Introduction
Extensive experiments demonstrate that UASR-LLM consistently outperforms state-of-the-art baselines across ASR, VSR, and AVSR when trained with comparable amounts of audio-visual data. On the LRS3 benchmark, our model achieves word error rates of 20.9%, 0.84%, and 0.69% for VSR, ASR, and AVSR, respectively. Notably, under noisy conditions, UASR-LLM exhibits strong robustness, yielding substantial WER reductions over competing methods. We further show that the proposed framework generalizes well across different choices of speech foundation models and large language models, with all configurations surpassing baseline performance. Finally, ablation studies validate the effectiveness of our training strategy and confirm the advantages of unified speech recognition within a single model architecture.
II Related works
II-A Speech foundation models
@@
Recent years have witnessed the emergence of various speech foundation models, which can be broadly categorized into two paradigms: self-supervised [24, 27, 28] and supervised approaches [5, 29]. Among SSL-based speech foundation models, HuBERT [27] utilizes k-means clustering to generate pseudo-class labels for model training using a mask prediction loss, with iterative refinement through subsequent clustering and mask prediction steps. WavLM [27] improves HuBERT by using a more diverse pretraining dataset and performs speech denoising modeling during pretraining. On the other hand, Whisper family [5] exemplifies supervised learning-based speech foundation model. Trained on large-scale multilingual and multitask labeled data, it demonstrates robust performance in ASR and speech translation tasks.
II-B Visual representation learning
@@
Visual representation learning is crucial for the performance of VSR and AVSR tasks, which can be broadly categorized into supervised and self-supervised approaches. Supervised methods leverage labeled [30] or weakly labeled audio-visual speech data [31] to optimize representation extraction and text decoding in an end-to-end manner through label prediction. Several auxiliary tasks have been proposed to enhance visual encoder learning, including articulation feature prediction [32] and distillation from acoustic models [32, 33, 34]. Self-supervised methods utilize unlabeled audio-visual data with pretext tasks such as masked prediction loss functions. Representative approaches include AV-HuBERT [15, 35], RAVEn [18, 16], AV2vec [36], and AV-data2vec [37]. Zhang et al. [23] proposed a joint audio-visual representation learning method that leverages cross-modal knowledge distillation from pretrained SFMs. Here, we further leverage SFMs directly for processing audio-visual data, which can be viewed as an extension of this study.
II-C Speech recognition with LLMs
@@
Large language models (LLMs) have attracted significant attention for speech recognition applications, with research broadly falling into two categories: multimodal LLMs for universal AI assistants [38, 39] and LLMs for enhancing traditional speech recognition tasks [40, 41, 42]. Our work belongs to the latter category. Recent approaches for integrating LLMs with speech recognition typically employ lightweight adapters to connect frozen speech encoders with LLMs. Representative works include joint speech-language models with small adapters [40] and Q-Former-based connections [42]. Some studies explore discrete speech representations as LLM inputs [38, 43]. LLMs have also been adopted for visual and audio-visual speech recognition [44, 45, 46]. VSP-LLM [44] focuses on visual speech recognition, while Llama-AVSR [46] uses separate audio and visual encoders with temporal concatenation as LLM input. MMS-Llama [47] employs audio-visual fusion within encoders and Q-former adapters for improved efficiency. Our work differs by proposing a unified multimodal speech recognition approach that processes both audio and visual modalities within a single model architecture, rather than targeting specific modalities separately.
III Proposed method
@@
Our UASR-LLM method comprises two main components: an audio-visual adapted speech foundation model (SFM-AV) that processes both audio and visual modalities and encodes them into a unified representational space, and a large language model that decodes the encoded representations into text. The SFM-AV consists of a visual encoder and a speech foundation model enhanced with visual injection modules. Our model architecture is illustrated in Figure 1. The training of our model follows a two-stage process: in the first pretraining stage, the SFM is adapted to encode both audio and visual information into a unified hidden space; in the second finetuning stage, the adapted SFM is connected to the LLM to enable unified multimodal speech recognition. The detailed model architecture and training strategy are described in the following sections.
III-A Model structure
III-A1 Visual encoder
@@
Let lip video be , where and represent the temporal, height, width, and channel dimensions, respectively. The visual encoder processes the visual input into hidden representations: , where is the feature dimension of the encoded representations. In this work, we adopt the visual component of the pretrained DistillAV model [23] as our visual encoder. DistillAV is designed to distill knowledge from speech foundation models (SFMs) for joint audio-visual representation learning. Unlike SFMs, which are trained on large-scale speech-only corpora, DistillAV is trained on relatively smaller audio-visual datasets. As a result, its performance in audio-only tasks is inferior to that of its teacher models, as shown in previous ASR evaluations [23]. However, DistillAV exhibits strong visual modeling capabilities and consistently outperforms state-of-the-art lipreading models such as AV-HuBERT [15] and AV2vec [36]. Therefore, we use the pretrained DistillAV model to extract representations from lip video inputs.
III-A2 Visual injection module
@@ III-A3 Prompting the LLM for USR
III-B Training strategy
III-B1 Visual injection pretraining
@@
where represents instance normalization. This multi-layer aggregation captures rich hierarchical features, which follows previous study [23].
@@
where represents learnable projection parameters.
III-B2 Speech recognition finetuning
@@
IV Experiments and results
IV-A Implementation details
@@
Model configurations. Our experiments employed a default configuration consisting of three main components: the DistillAV model [23] as the visual encoder, WavLM-large [4] (317M parameters) as the speech foundation model (SFM), and Qwen 2.5-7B555https://huggingface.co/Qwen/Qwen2.5-7B as the large language model. The visual encoder offered two variants: Base (103M parameters) and Large (325M parameters). Following previous studies [15, 23], we denote the proposed method as UASR-LLM-B and UASR-LLM-L accordingly. The visual injection modules were configured with 8 heads, 512 hidden dimensions, and 256 dimensions for both cross-attention and feed-forward networks, totaling approximately 58M parameters. The two-layer feed-forward network adapter that connects the SFM and LLMs contains around 10M parameters. LoRA weights were configured with a rank of 16, resulting in 12M total parameters. To demonstrate the generalization capability of our method, we explored additional speech foundation models: the open-source Whisper-medium [5] and an internal Conformer-based ASR model (Conformer-Speech), both containing approximately 300M parameters comparable to WavLM-large. We also evaluated ChatGLM v3-6B666https://huggingface.co/THUDM/chatglm3-6b as an alternative large language model.
The WavLM model generates representations at 50 fps, double the frame rate of visual representations. We set relative positional encoding according to the real-time interval between audio and visual frames using sinusoidal positional embedding. The parameter in Equation 4 was set to 8 for WavLM-based SFM and 1 for Whisper and Conformer-Speech SFM, following the previous study [23].
@@
IV-B Comparison with state-of-the-art baselines
| Methods | Unlab. | Lab. | Shared | VSR | ASR | AVSR | ASR (Noisy) | AVSR (Noisy) |
@@ IV-B Comparison with state-of-the-art baselines
| Methods | Unlab. | Lab. | Shared | VSR | ASR | AVSR | ASR (Noisy) | AVSR (Noisy) |
@@
Table I and Table II present experimental results for models finetuned with low-resource (30h) and high-resource (433h or 1759h) labeled data, respectively. The unlabeled data refers to audio-visual data used for pretraining the DistillAV-based visual encoder and visual injection pretraining, while the labeled data denotes audio-visual data employed for speech recognition finetuning.
As shown in Table I, our method outperforms most baselines across VSR, ASR, and AVSR tasks under both clean and noisy conditions. Although underperforming Llama-AVSR in ASR, our approach provides a unified framework effective across all three tasks. The superior VSR performance particularly demonstrates the effectiveness of ”reprogramming” WavLM for lipreading with LLMs, consistent with previous studies [20]. With 433h labeled data, as shown in Table II, our method achieves significant accuracy improvements. Compared to previous studies using comparable audio-visual data, our approach demonstrates superior performance across all tasks under both clean and noisy conditions. With 1759h labeled data, our method achieves WERs of 20.9%, 0.84%, and 0.69% on the clean LRS3 test set, rivaling or surpassing state-of-the-art methods. However, our VSR results still lag behind LP-Conformer [53], which uses approximately 100k hours of audio-visual data.
@@
IV-C AVSR results on noisy test sets
To evaluate robustness under noisy conditions, we constructed four test sets by mixing the clean audio from the LRS3 test set with different noise types: speech, babble, natural, and music, as described in Section IV-A. We compared our approach against multiple baselines, including two audio-only SFM models (WavLM and Conformer-Speech) and three AVSR models (AV-HuBERT [15], AV2vec-MLM [36], and DistillAV [23]). All audio-only models were finetuned on the LRS3 dataset. The experiments utilized 433h data for pretraining, with all models adopting the Base configuration. We used either 30h or 433h of LRS3 data for finetuning. The experimental results are summarized in Figure 4.
Our proposed UASR-LLM method (red line) consistently outperforms all baselines across different noise types and SNR levels, with particularly pronounced advantages in severe noise conditions below 0 dB. Among the four noise conditions, babble noise proves most challenging, with all methods exhibiting higher WER at low SNRs, particularly below 0 dB. However, the proposed approach maintains substantially lower error rates than competing methods even in these challenging low-SNR scenarios. The results highlight the effectiveness of incorporating visual information, as audio-only models (WavLM and Conformer-Speech) degrade rapidly in noisy conditions compared to AVSR approaches.
| Labeled data | LLM type | SFM type | VSR | ASR | AVSR | ASR (Noisy) | AVSR (Noisy) |
| DistillAV [23] | |||||||